OpenStudy (anonymous):

If the similarity ratio of two similar solids is 3 : 14, what is the ratio of their corresponding areas? What is the ratio of their corresponding volumes? A. The ratio of their corresponding areas is 27 : 2744. The ratio of their corresponding volumes is 9 : 196. B. The ratio of their corresponding areas is 9 : 196. The ratio of their corresponding volumes is 27 : 2744. C. The ratio of their corresponding areas is 6 : 28. The ratio of their corresponding volumes is 9 : 42. D. The ratio of their corresponding areas is 3 : 196. The ratio of their corresponding volumes is 3 : 2744.

OpenStudy (nubeer):

\[\frac{ A1 }{ A2 } = \left( \frac{ L1 }{ L2 } \right)^{2}\]